Find the inverse Laplace trans form (using table)
Answer #1

| S之 let F() = = ) | F(O) : +4 - )() Fit) = + + - 3 GL 十一 Inverse Laplace transform [roj = 1 + [3] [4] ( 5 21 ] -- - + [ | -}S- sta -35 WWW 2 Let Fly Tree de fls) = 5 + 235 14(coshat) = - la(sin hat) = a 22 FLD = { 13 tahansa 31-05) Q2 BIT √35 Taking- S + T - 3 (5-3 ) 3 Fis) = Taking Laplace 5 + 3(5-3) Inverse → erse 34, 7 t 5 sta 5²-35 dlm ال) ANSWER. Thats why I Note:-